# arXiv:2603.01290v1[cs.AI] 01 Mar 2026

### v1

@@ Abstract
The 2026 Formula 1 technical regulations introduce a fundamental change to energy strategy: under a 50/50 internal combustion engine / battery power split with unlimited regeneration and a driver-controlled Override Mode (hereafter: Override Mode; abbreviated MOM throughout), the optimal energy deployment policy depends not only on a driver’s own state but on the hidden state of rival cars. This creates a Partially Observable Stochastic Game that cannot be solved by single-agent optimisation methods.
We present a tractable two-layer inference and decision framework. The first layer is a 30-state Hidden Markov Model (HMM) that infers a probability distribution over each rival’s ERS charge level, Override Mode status, and tyre degradation state from five publicly observable telemetry signals. The second layer is a Deep Q-Network (DQN) policy that takes the HMM belief state as input and selects between energy deployment strategies. We formally characterise the counter-harvest trap — a deceptive strategy in which a car deliberately suppresses observable deployment signals to induce a rival into a failed attack — and show that detecting it requires belief-state inference rather than reactive threshold rules.
@@ 1 Introduction
The 2026 regulations change this (Fédération Internationale de l’Automobile, 2025). The 50/50 power split between internal combustion engine and MGU-K battery places the battery under permanent demand, making energy state a first-order strategic variable every sector. Override Mode is a proximity-gated energy award: when a driver is within one second of the car ahead at the designated detection point, they earn an extra 0.5MJ of deployable energy on the following lap (Fédération Internationale de l’Automobile, 2025). This replaces the function of DRS as a proximity-based overtaking aid — the mechanism changes from a drag-reduction wing to an energy award, but the strategic role is similar. Active Aero replaces the hardware of DRS entirely: the adjustable rear wing is removed and replaced with a driver-controlled binary aero system available only in designated activation zones. Straight-line mode provides mechanical speed independent of ERS deployment, which is the physical basis for the counter-harvest trap. The optimal burn-or-harvest decision in any sector is now a function of the joint state of both cars, not just one’s own.
@@
    Problem formalisation. We formally model 2026 F1 energy management as a Partially Observable Stochastic Game (POSG) (Hansen et al., 2004) and define a tractable POMDP approximation for the single-agent case in which rivals are treated as stationary processes. We precisely characterise the counter-harvest trap — a deceptive equilibrium strategy whose existence motivates the full multi-agent treatment of Papers 2 and 3 in this series.
@@
    Rival state inference. We propose a 30-state HMM over rival ERS charge level, Override Mode status, and tyre degradation, initialised from an analytically-derived emission matrix and calibrated via Baum-Welch EM on 2026 race telemetry. We identify and document that Override Mode status is inferred primarily through transition dynamics and partially-observable proximity conditions, and that tyre emission signal is concentrated in the cliff degradation state.
@@
    Decision policy. We propose and fully specify a DQN policy trained on the HMM belief state, and characterise its expected advantages over observable-only threshold baselines under the synthetic evaluation regime. Empirical policy comparison against real-data baselines is deferred to post-Melbourne validation.
Scope and limitations.
This paper treats rivals as stationary processes — their policies do not adapt to our inference. This assumption is violated when rivals reason about being observed, leading to the counter-harvest trap and non-stationary equilibria studied in Paper 2 in this series. We state this limitation at the outset because it is not a defect to be hidden: it is precisely the gap between Paper 1 and Paper 2. The POMDP solution presented here is valuable despite this limitation: it is the correct baseline against which the game-theoretic extension must demonstrate improvement.
@@ 2 Background
Prior computational work on F1 strategy has focused on pitstop timing as an optimisation problem (Heilmeier et al., 2020) and on reinforcement learning for race simulation (Betz et al., 2022). These formulations treat energy management as a deterministic sub-problem with known rival actions. Betz et al. (2022) model ERS deployment in hybrid powertrains but in a single-agent setting without rival inference.
The closest structural analogue to the counter-harvest trap in prior work is the Formula E Attack Mode: teams earn a temporary power boost by driving through a designated off-line zone, creating a strategic asymmetry where a car can accept a short-term pace penalty to gain a power advantage (Heilmeier et al., 2020). The key difference from the 2026 counter-harvest trap is that Attack Mode is opt-in from a fixed physical location, publicly visible to all teams, and does not interact with a rival’s inference model — the power boost is simply announced. The 2026 Active Aero system creates a strictly harder version of this problem: the equivalent of Attack Mode can be executed covertly at any activation zone, at a time of the driver’s choosing, in a way that mimics harvesting to an observer without a belief-state model.
To our knowledge no prior work models 2026-specification Override Mode or Active Aero under adversarial conditions.
@@
[…]
@@
With 84,000 emission parameters and $\sim$3,480 sector observations per rival per race, the model is unidentifiable from a single race weekend — Baum-Welch estimates from Race 1 are Dirichlet-regularised guesses, not reliable posteriors. By Race 4, cumulative observations reach $\sim$13,920, giving a parameter-to-data ratio of approximately 6:1. With parameter tying by (ERS, tyre) cluster reducing to 15 effective emission rows, the ratio improves to $\sim$1.5:1 at Race 4, at which point the EM estimates become meaningful. The model should be treated as empirically validated only from Race 4 onwards; Melbourne results are explicitly preliminary under this interpretation.
@@
DQN (Mnih et al., 2015) with double Q-learning (van Hasselt et al., 2016) and experience replay (Lin, 1992) provides the policy learning component. We use a deliberately shallow network (three hidden layers, 256–256–128 units) because the belief state does the representational heavy lifting and the training data volume is limited to $\sim$3,480 sector observations per car per season.
@@ 3.1 The 2026 Regulatory Context
The 2026 regulations introduce three strategically relevant changes (Fédération Internationale de l’Automobile, 2025):
@@
The total power output is split equally between the internal combustion engine and the MGU-K battery system. Unlike prior hybrid regulations, the battery is under continuous demand, making ERS charge level a first-order strategic variable at every sector rather than only under overtaking.
@@
A proximity-gated energy award replacing DRS entirely. When a driver is within one second of the car ahead at the lap’s detection point (nominally the final corner), they earn an extra 0.5MJ of deployable energy on the following lap (Fédération Internationale de l’Automobile, 2025). When deployed, Override Mode produces a distinctive speed signature detectable in the speed trap telemetry. Unlike a push-to-pass system, Override Mode is earned through proximity rather than free to use at any time. The strategic value of a deployment depends on whether the rival has sufficient ERS charge to defend against the attack.
@@
Replaces DRS entirely. Driver-controlled, available only in designated activation zones, with two fixed positions (cornering and straight-line). Straight-line mode in an activation zone produces a speed boost detectable as a positive $\Delta\text{speed\_trap}$ increment above the ERS-explained baseline. Crucially, a car can run Active Aero in straight-line mode while harvesting (ERS==L) — the mechanical downforce reduction provides speed independent of battery deployment. This is the physical basis for the counter-harvest trap (Section 3.3).
@@
Pre-season analysis of the 2026 cars reveals that the recharge-per-lap ratio varies from 1.0 at energy-scarce circuits (Australia, Italy, Saudi Arabia) to 2.2 at regen-rich circuits (Azerbaijan, Singapore). [^1] The recharge-per-lap ratios, the 16-second super clipping figure for Australia, and the 250 kW recharge rate are community-derived engineering estimates from pre-season analysis circulating in February 2026. The authors were unable to identify a primary attributable source for these figures; they are not FIA-certified. All downstream claims that depend on them (Sections 5 and 8) should be treated as provisional until confirmed by Race 1 telemetry. At Melbourne specifically, cars require approximately 16 additional seconds of super clipping per lap at 250 kW to approach the energy ceiling imposed by the 50/50 power split. The MGU-K battery capacity is approximately 4 MJ (at 350 kW peak discharge rate, the battery depletes in $\sim$11.4 s); at 250 kW over 16 s this represents a full extra battery cycle beyond what natural regen provides. This transforms super clipping from a discretionary tool into a near-mandatory operating mode at certain circuits, with direct consequences for the counter-harvest trap’s observability discussed in Section 3.3.
@@ 3.2 The POMDP Model
The fully observable component $s^{o}_{t}$ includes tyre compound and age (FIA-published), lap number, sector, circuit position (activation zone or not), and own-car ERS level, Override Mode status, Active Aero position, and gap to adjacent cars. This information is available to the ego car without inference.
@@
$e$ $\in\{H,M,L\}\quad\text{(ERS charge level)}$ (1)
$m$ $\in\{\text{available},\text{spent}\}\quad\text{(Override Mode status)}$ (2)
$\tau$ $\in\{\text{new},\text{light},\text{moderate},\text{heavy},\text{cliff}\}\quad\text{(tyre degradation)}$ (3)
giving $|\mathcal{X}|=3\times 2\times 5=30$ states.
@@
Fuel load is excluded from the hidden state for two independent reasons: (i) all cars start with the same regulated load and deplete it at approximately the same rate per lap, making fuel state deterministic from the observable lap number; and (ii) all five observables are measured as deviations from a per-driver rolling 5-lap baseline, which absorbs the smooth monotone pace improvement from fuel burn. We verified this directly: the fuel parameter was accepted by our emission model but produced zero differentiation across any observable computation.
@@
$\mathcal{A}=\{\text{burn},\text{harvest}\}$. Burn denotes ERS deployment above the per-sector baseline rate. Harvest denotes deployment below baseline, recovering charge through regeneration.
@@
$o_{t}=(\Delta v_{\text{trap}},\Delta t_{\text{sector}},\Delta b_{\text{brake}},\sigma^{2}_{\text{speed}},z_{\text{aero}})\in\Omega$ where each component is a deviation from the rival’s rolling 5-lap baseline.
The five observables are: (i) $\Delta v_{\text{trap}}$: speed at the speed trap minus rival baseline (km/h); (ii) $\Delta t_{\text{sector}}$: sector time minus rival baseline (seconds); (iii) $\Delta b_{\text{brake}}$: distance from apex at which braking begins, minus rival baseline (metres); (iv) $\sigma^{2}_{\text{speed}}$: detrended variance of rival’s speed trace within the sector — elevated during inconsistent deployment patterns; (v) $z_{\text{aero}}\in\{0,1\}$: whether Active Aero is deployed in straight-line mode (observable in activation zones from speed signature).
@@
$b_{t}\in\Delta^{30}$ is the probability vector over hidden states, maintained by the HMM forward algorithm:
@@
$R(s_{t},a_{t})=\mathbb{E}\left[\sum_{k=1}^{H}\gamma^{k}\cdot\Delta\text{pos}_{t+k}\;\middle|\;s_{t},a_{t}\right]$ (5)
where $\Delta\text{pos}$ is the change in race position, $H=5$ (planning horizon covering approximately two ERS cycles), and $\gamma=0.95$.
@@ 3.3 The Counter-Harvest Trap
A deceptive strategy for a car ahead (Car B) against a following car (Car A) in which Car B simultaneously: (i) runs ERS in harvest mode (ERS==L, conserving charge); (ii) deploys Active Aero in straight-line mode within an activation zone (maintaining straight-line speed despite harvesting); and (iii) allows the observable telemetry signals to indicate harvesting. Car A’s inference model identifies the ERS==L signature as an attack opportunity and burns energy. Car B then activates full deployment (including Override Mode), defending the position easily because it has the full energy reserve that Car A just spent.
The trigger signature exploits a key 2026 regulatory asymmetry: lift-and-coast disables active aero, but super clipping does not. Super clipping is the automated redirection of ICE power to the battery at full throttle; the car slows slightly (power diverted from the rear wheels) but active aero remains deployed (Fédération Internationale de l’Automobile, 2025).
The counter-harvest trap therefore has a precise observable signature: $z_{\text{aero}}=1$ (Active Aero deployed, consistent with full throttle) combined with $\Delta v_{\text{trap}}$ far below what committed Active Aero deployment with full energy would produce. A car genuinely attacking with Active Aero deployed produces $\Delta v_{\text{trap}}\approx+8$ km/h; the trap car produces only $\approx+2$ km/h, indistinguishable from ERS==M without Active Aero to an observer who does not track $z_{\text{aero}}$. A rival observing only $\Delta v_{\text{trap}}$ sees +2+2 km/h and interprets this as moderate energy state and no attack opportunity; in reality the battery is filling and the Active Aero is concealing it. A model without Active Aero as an explicit observable cannot distinguish super clipping from genuine ERS==L without deception.
@@
The counter-harvest trap is not detectable by a single-observable threshold policy on $\Delta v_{\text{trap}}$ alone.
@@
A car running ERS==H without Active Aero produces $\Delta v_{\text{trap}}\approx+4$ km/h. A car running ERS==L with Active Aero in an activation zone produces $\Delta v_{\text{trap}}\approx-2+4=+2$ km/h. These distributions overlap under the emission model (Section 4). No scalar threshold on $\Delta v_{\text{trap}}$ can reliably separate them: the trap car’s +2+2 km/h is indistinguishable from ERS==M (+1+1 km/h) without additional information. The joint signal $(\Delta v_{\text{trap}},z_{\text{aero}})$ provides separation: $z_{\text{aero}}=1$ with $\Delta v_{\text{trap}}\ll+8$ km/h (specifically, $\Delta v_{\text{trap}}$ in the range $[0,+3]$ km/h) identifies the trap condition with probability $\geq 0.80$ under the emission model, because a genuinely attacking car with Active Aero would produce $\approx+8$ km/h. ∎
@@
The diagnostic power of the $z_{\text{aero}}=1$ with depressed $\Delta v_{\text{trap}}$ signature depends on that combination being unusual — its rarity is what gives it discriminatory value. This assumption holds at regen-rich circuits but breaks down at energy-scarce ones.
At circuits where super clipping is mandatory for approximately 16 seconds per lap (Section 3.3), the active super-clipping speed penalty is large enough that $\Delta v_{\text{trap}}<0$ during straight-line phases is the expected modal state of every car on the grid, regardless of strategic intent. In this regime, even the corrected trap signature — $z_{\text{aero}}=1$ with depressed $\Delta v_{\text{trap}}$ — becomes ambient: every car running Active Aero while super-clipping produces exactly this pattern. The combination is therefore ambient rather than diagnostic at circuits like Australia.
The rolling 5-lap baseline provides a partial rescue. Since all cars are super clipping in the same track sections, the baseline adapts and treats mandatory super clipping as the new normal. A car that has conserved energy and avoided super clipping will still produce a positive $\Delta v_{\text{trap}}$ relative to the super-clipping field. The baseline-delta design therefore partially preserves signal even under universal super clipping, but the margin for detecting deliberate deception narrows considerably relative to regen-rich circuits.
The implication is that trap detection confidence is circuit-dependent, correlated with the recharge-per-lap ratio. At Azerbaijan (ratio 2.2), super clipping is rare and the signature retains strong diagnostic value. At Australia (ratio 1.0), the same signature is the default operating condition and the trap becomes substantially harder to distinguish from normal racing. This generates a testable prediction: trap detection recall should be lower at Australia than the synthetic baseline of 95.7%, and should improve at higher-regen circuits later in the season.
Critically, this circuit dependence does not invalidate the framework — it refines the conditions under which the counter-harvest trap is a viable strategic tool. We conjecture that a rational trap-setter prefers high-regen circuits, where their deception operates against a cleaner background signal and the $z_{\text{aero}}=1$ with depressed $\Delta v_{\text{trap}}$ combination retains diagnostic rarity. Whether this conjecture holds depends on a non-trivial game-theoretic question: at high-regen circuits the defender’s HMM also operates with cleaner signal and tracks ERS state more accurately, potentially making the trap harder to execute as well as easier to detect. Resolving this requires the multi-agent analysis of Paper 2; we present the circuit-dependence observation as empirical motivation rather than a settled strategic prediction. At low-regen circuits, the signal confusion is symmetric: both attacker and defender operate with degraded information, and the advantage shifts to whichever car manages its energy state more reliably over a stint.

### v2

# Opponent State Inference Under Partial Observability: An HMM–POMDP Framework for 2026 Formula 1 Energy Strategy Kalliopi Kleisarchaki

@@ Abstract
The 2026 Formula 1 technical regulations introduce a fundamental change to energy strategy: under a 50/50 internal combustion engine / battery power split with unlimited regeneration and a driver-controlled Override Mode (hereafter mom), the optimal energy deployment policy depends not only on a driver’s own state but on the hidden state of rival cars. This creates a Partially Observable Stochastic Game that cannot be solved by single-agent optimisation methods.
We present a tractable two-layer inference and decision framework. The first layer is a 40-state Hidden Markov Model (hmm) that infers a probability distribution over each rival’s ers charge level—now decomposed into four modes ($H$, $M$, $L_{\mathrm{harvest}}$, $L_{\mathrm{derate}}$)—Override Mode status, and tyre degradation state from six publicly observable telemetry signals. The second layer is a Deep Q-Network (dqn) policy that takes the hmm belief state as input and selects between energy deployment strategies. We formally characterise the counter-harvest trap—a deceptive strategy in which a car deliberately suppresses observable deployment signals to induce a rival into a failed attack—and show that detecting it requires belief-state inference over both ERS level and the harvest/derate sub-mode.
The $L_{\mathrm{harvest}}$$L_{\mathrm{derate}}$ split is the central architectural advance of v2: both sub-modes produce depressed $\Delta v_{\mathrm{trap}}$ and can co-occur with $z_{\mathrm{aero}}{}=1$ in an activation zone, but they carry opposite strategic implications. $L_{\mathrm{harvest}}$ means the rival is building a hidden reserve—the trap condition. $L_{\mathrm{derate}}$ means the rival is at their physical limit—a genuine attack opportunity. The sixth observable $\delta_{\mathrm{throttle}}$ (super-clipping duration fraction, added in v1.5) provides clean separation between these sub-modes; v2 elevates this separation to the state level rather than handling it as a mixed emission.
@@ 1 Introduction
The 2026 regulations change this (FIA, 2025). The 50/50 power split between internal combustion engine and MGU-K battery places the battery under permanent demand, making ers charge level a first-order strategic variable every sector. Override Mode is a proximity-gated energy award: when a driver is within one second of the car ahead at the designated detection point, they earn an extra 0.5 MJ of deployable energy on the following lap. Active Aero replaces DRS entirely: the adjustable rear wing is removed and replaced with a driver-controlled binary aero system available only in designated activation zones. Crucially, a car can run Active Aero in straight-line mode while harvesting—the mechanical downforce reduction provides speed independent of battery deployment. This is the physical basis for the counter-harvest trap (Section 3.3).
The optimal burn-or-harvest decision in any sector is now a function of the joint state of both cars, not just one’s own.
Contributions
@@
    Problem formalisation. We formally model 2026 F1 energy management as a Partially Observable Stochastic Game (posg) and define a tractable pomdp approximation for the single-agent case. We precisely characterise the counter-harvest trap as a deceptive equilibrium strategy.
@@
    Rival state inference. We propose a 40-state hmm over rival ers charge level (four modes: $H$, $M$, $L_{\mathrm{harvest}}$, $L_{\mathrm{derate}}$), Override Mode status, and tyre degradation, initialised from an analytically-derived emission matrix and calibrated via Baum-Welch EM on 2026 race telemetry. The $L_{\mathrm{harvest}}$$L_{\mathrm{derate}}$ decomposition is the architectural advance over v1.5: it eliminates the ambiguity that v1.5 handled only at the emission level via a mixed prior.
@@
    Decision policy. We propose and fully specify a dqn policy trained on the 40-dimensional hmm belief state, and characterise its advantages over observable-only threshold baselines.
Scope and limitations
This paper treats rivals as stationary processes—their policies do not adapt to our inference. This assumption is violated when rivals reason about being observed, leading to the counter-harvest trap and non-stationary equilibria studied in Kleisarchaki (2026b). We state this limitation at the outset because it is not a defect to be hidden: it is precisely the gap between this paper and its game-theoretic extension. The pomdp solution presented here is the correct baseline against which that extension must demonstrate improvement.
@@ 2 Background
Prior computational work on F1 strategy has focused on pit-stop timing as an optimisation problem (Heilmeier et al., 2020) and on reinforcement learning for race simulation (Betz et al., 2022). These formulations treat energy management as a deterministic sub-problem with known rival actions.
The closest structural analogue to the counter-harvest trap in prior work is the Formula E Attack Mode (Heilmeier et al., 2020): teams earn a temporary power boost by driving through a designated off-line zone. The key difference is that Attack Mode is opt-in from a fixed physical location, publicly visible to all teams, and does not interact with a rival’s inference model. The 2026 Active Aero system creates a strictly harder version: the equivalent of Attack Mode can be executed covertly at any activation zone, at a time of the driver’s choosing, in a way that mimics harvesting to an observer without a belief-state model.
@@
Why discrete HMM over particle filter.
A particle filter could in principle handle continuous ers charge and relax the conditional independence assumption. We use a discrete hmm for two reasons. First, Baum-Welch provides closed-form maximum-likelihood parameter estimation from unlabelled sequences; particle filters require approximate inference schemes (e.g. SMC-EM) that are substantially less stable at the small data volumes available from a single race weekend. Second, the four-bin ers discretisation is physically motivated: the strategic decision is whether a rival can defend ($\textsc{ers}\geq M$), cannot defend ($L_{\mathrm{derate}}$: physically depleted), or is building a trap ($L_{\mathrm{harvest}}$: deliberately conserving); a continuous representation would require decision boundaries to be learned from data rather than specified from domain knowledge.
@@
With 40 states and six observables discretised to $8\times 7\times 5\times 5\times 2\times 6=16{,}800$ joint bins, the raw emission matrix has $40\times 16{,}800=672{,}000$ parameters. With parameter tying by (ERS, tyre) cluster (Section 4) reducing to 20 effective emission rows, and exploiting the conditional independence structure, the effective degrees of freedom per emission row are 8+7+5+5+2+65=288{+}7{+}5{+}5{+}2{+}6{-}5=28, giving $20\times 28=560$ effective parameters. Against $\sim$3,480 sector observations per rival per race, the model reaches a meaningful parameter-to-data ratio of $\sim$1.6:1 by Race 4 ($\sim$13,920 observations). The model should be treated as empirically validated only from Race 4 onwards; Melbourne results are explicitly preliminary.
@@
DQN (Mnih et al., 2015) with double Q-learning (van Hasselt et al., 2016) and experience replay (Lin, 1992) provides the policy learning component. We use a deliberately shallow network (three hidden layers, 256–256–128 units) because the belief state does the representational heavy lifting and the training data volume is limited.
@@ 3.1 The 2026 Regulatory Context
The 2026 regulations introduce three strategically relevant changes (FIA, 2025).
@@
Total power output is split equally between ICE and MGU-K battery. The battery is under continuous demand, making ers charge level a first-order strategic variable at every sector.
@@
A proximity-gated energy award replacing DRS entirely. When a driver is within one second of the car ahead at the lap’s detection point (nominally the final corner), they earn an extra 0.5 MJ of deployable energy on the following lap. When deployed, Override Mode produces a distinctive speed signature detectable in speed-trap telemetry.
@@
Replaces DRS entirely. Driver-controlled, available only in designated activation zones, with two fixed positions (cornering and straight-line). Straight-line mode produces a speed boost detectable as a positive $\Delta v_{\mathrm{trap}}$ increment above the ers-explained baseline. Crucially, a car can run Active Aero in straight-line mode while super-clipping ($L_{\mathrm{derate}}$)—the mechanical downforce reduction provides speed independent of battery deployment. This is the physical basis for the counter-harvest trap.
@@
[…]
@@ 3.2 The POMDP Model
The fully observable component $s^{o}_{t}$ includes tyre compound and age (FIA-published), lap number, sector, circuit position (activation zone or not), and own-car ers level, Override Mode status, Active Aero position, and gap to adjacent cars.
@@
$e$ $\in\{H,\;M,\;L_{\mathrm{harvest}},\;L_{\mathrm{derate}}\}$ (ers charge level, 4 modes) (1)
$m$ $\in\{\mathtt{available},\;\mathtt{spent}\}$ (Override Mode status) (2)
$\tau$ $\in\{\mathtt{new},\;\mathtt{light},\;\mathtt{moderate},\;\mathtt{heavy},\;\mathtt{cliff}\}$ (tyre degradation) (3)
giving $|\mathcal{X}|=4\times 2\times 5=\mathbf{40}$ states.
The $L_{\mathrm{harvest}}$ / $L_{\mathrm{derate}}$ decomposition.
This is the central architectural advance of v2 over v1.5. Both sub-modes produce depressed $\Delta v_{\mathrm{trap}}$ and may co-occur with $z_{\mathrm{aero}}=1$ in an activation zone, but they have opposite strategic implications:
  • $L_{\mathrm{harvest}}$: the driver is actively managing throttle below 100%, building a hidden energy reserve. This is the trap condition. $\delta_{\mathrm{throttle}}\approx 0.05$$0.10$ (low; managed throttle, not full demand).
  • $L_{\mathrm{derate}}$: the driver demands full power ($\geq$98% throttle) but the MGU-K battery has reached its SOC ceiling and cannot deliver further electrical output. The car slows not by choice but by physics. $\delta_{\mathrm{throttle}}\approx 0.40$$0.60$ (high; full throttle with below-baseline speed).
In v1.5, both sub-modes shared the $L$ state, and the $\delta_{\mathrm{throttle}}$ observable was handled as a mixed emission with a circuit-dependent prior ($\mu=0.40$ for Melbourne at recharge ratio 1.0). This worked, but produced an ambiguous posterior over $L$: the hmm could not separately report $P(L_{\mathrm{harvest}})$ and $P(L_{\mathrm{derate}})$, which are the quantities the dqn policy actually needs. In v2, the posterior directly provides both quantities as separate belief dimensions.
@@
Fuel load is excluded from the hidden state for two independent reasons: (i) all cars start with the same regulated load and deplete it at approximately the same rate, making fuel state deterministic from the observable lap number; and (ii) all six observables are measured as deviations from a per-driver rolling 5-lap baseline, which absorbs the smooth monotone pace improvement from fuel burn.
@@
$A=\{\mathtt{burn},\;\mathtt{harvest}\}$.
@@
$o_{t}=(\Delta v_{\mathrm{trap}},\Delta t_{\mathrm{sector}},\Delta b_{\mathrm{brake}},\sigma^{2}_{\mathrm{speed}},z_{\mathrm{aero}},\delta_{\mathrm{throttle}})\in\Omega$, where each component is a deviation from the rival’s rolling 5-lap baseline, except $z_{\mathrm{aero}}\in\{0,1\}$ and $\delta_{\mathrm{throttle}}\in[0,1]$.
The six observables are: (i) $\Delta v_{\mathrm{trap}}$: speed at the speed trap minus rival baseline (km/h); (ii) $\Delta t_{\mathrm{sector}}$: sector time minus rival baseline (seconds); (iii) $\Delta b_{\mathrm{brake}}$: distance from apex at which braking begins, minus rival baseline (metres); (iv) $\sigma^{2}_{\mathrm{speed}}$: detrended variance of rival’s speed trace within the sector; (v) $z_{\mathrm{aero}}\in\{0,1\}$: Active Aero deployment in straight-line mode (observable in activation zones from speed signature); (vi) $\delta_{\mathrm{throttle}}\in[0,1]$: fraction of straight-line telemetry samples within the sector where throttle $\geq 98\%$ and speed is simultaneously below the driver’s 5-lap rolling baseline—the mechanical signature of ers derating, added in v1.5 and now elevated to full state-level discriminator in v2.
@@
$b_{t}\in\Delta^{40}$ is the probability vector over hidden states, maintained by the hmm forward algorithm:
@@
$R(s_{t},a_{t})=\mathbb{E}\!\left[\sum_{k=1}^{H}\gamma^{k}\cdot\Delta\mathrm{pos}_{t+k}\,\middle|\,s_{t},a_{t}\right]$ (5)
where $\Delta\mathrm{pos}$ is the change in race position, $H=5$ (planning horizon covering approximately two ers cycles), and $\gamma=0.95$.
@@ 3.3 The Counter-Harvest Trap
A deceptive strategy for a car ahead (Car B) against a following car (Car A) in which Car B simultaneously: (i) runs ers in $L_{\mathrm{harvest}}$ mode (conserving charge deliberately); (ii) deploys Active Aero in straight-line mode within an activation zone (maintaining straight-line speed despite harvesting); and (iii) allows the observable telemetry signals to indicate harvesting. Car A’s inference model identifies the $L$-state signature as an attack opportunity and burns energy. Car B then activates full deployment (including Override Mode), defending the position easily because it has the full energy reserve that Car A just spent.
The trigger signature exploits a key 2026 regulatory asymmetry: lift-and-coast disables Active Aero, but super-clipping does not. Super-clipping is the automated redirection of ICE power to the battery at full throttle; the car slows slightly (power diverted from rear wheels) but Active Aero remains deployed.
Why v2 resolves the trap ambiguity more cleanly than v1.5.
In v1.5, $L_{\mathrm{harvest}}$ and $L_{\mathrm{derate}}$ both mapped to the $L$ state. The $\delta_{\mathrm{throttle}}$ observable provided evidence to distinguish them, but the hmm’s posterior was expressed over $L$ as a single state—the policy had to infer the sub-mode from $\delta_{\mathrm{throttle}}$ directly rather than from a clean belief dimension. In v2, the posterior over $L_{\mathrm{harvest}}$ is a separate number from the posterior over $L_{\mathrm{derate}}$. The trap detection condition becomes simply:
$\text{trap condition}\;=\;P(L_{\mathrm{harvest}})>\theta_{\mathrm{trap}}\;\wedge\;z_{\mathrm{aero}}=1\;\wedge\;\text{in activation zone}$ (6)
This is directly actionable without post-hoc correction.
@@
The counter-harvest trap is not detectable by a single-observable threshold policy on $\Delta v_{\mathrm{trap}}$ alone.
@@
A car running $\textsc{ers}=H$ without Active Aero produces $\Delta v_{\mathrm{trap}}\approx+4$ km/h. A car in $L_{\mathrm{harvest}}$ with Active Aero in an activation zone produces $\Delta v_{\mathrm{trap}}\approx-2+4=+2$ km/h. A car in $L_{\mathrm{derate}}$ with Active Aero produces similarly $\Delta v_{\mathrm{trap}}\approx+2$ km/h. All three distributions overlap under the emission model. No scalar threshold on $\Delta v_{\mathrm{trap}}$ can reliably separate them. The joint signal $(\Delta v_{\mathrm{trap}},z_{\mathrm{aero}},\delta_{\mathrm{throttle}})$ provides separation: $z_{\mathrm{aero}}=1$ with $\Delta v_{\mathrm{trap}}\in[0,+3]$ km/h and low $\delta_{\mathrm{throttle}}$ (<0.15<0.15) identifies $L_{\mathrm{harvest}}$ with probability $\geq 0.85$ under the emission model. The same $(\Delta v_{\mathrm{trap}},z_{\mathrm{aero}})$ signature with high $\delta_{\mathrm{throttle}}$ (>0.35>0.35) identifies $L_{\mathrm{derate}}$—a genuine opportunity, not a trap. ∎
@@
At low-regen circuits (Melbourne, recharge ratio 1.0), super-clipping is mandatory for $\sim$16 s per lap, making $L_{\mathrm{derate}}$ the ambient state of most cars on the grid. The $z_{\mathrm{aero}}=1$ with depressed $\Delta v_{\mathrm{trap}}$ combination is therefore less diagnostic. The rolling 5-lap baseline provides a partial rescue—a car in $L_{\mathrm{harvest}}$ will still produce positive $\Delta v_{\mathrm{trap}}$ relative to the super-clipping field—but the margin narrows. The testable prediction is that trap-detection recall at Melbourne will be lower than the 96.3% synthetic baseline and will improve at higher-regen circuits.